OpenStudy (anonymous):

What is the simplified form of (1/x) - (2/x^2+x)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\frac{ 1 }{ x }-\frac{ 2 }{ x ^{2}+x } = \frac{ 1 }{ x }-\frac{ 2 }{ x(x+1) }=\frac{ x(x+1) -2x }{ x(x+1) } = \frac{ x ^{2}+x-2x }{ x(x+1) } = \frac{ x ^{2}-x }{ x(x+1) }=\frac{ x(x-1) }{ x(x+1) }= \frac{ x-1 }{ x+1 }\]

OpenStudy (anonymous):

\[\frac{ x(x-1) }{ x(x+1) } = \frac{ x-1 }{ x+1 }\]

OpenStudy (anonymous):

That is your final answer.